# Evaluating Cooling Center Coverage Using Persistent Homology of a Filtered Witness Complex

### v1

@@
Abstract
In light of the increase in frequency of extreme heat events, there is a critical need to develop tools to identify geographic locations that are at risk of heat-related mortality. This paper aims to identify locations by assessing holes in cooling-center coverage using persistent homology, a method from topological data analysis. Methods involving persistent homology have shown promising results in identifying holes in coverage of specific resources. We adapt these methods using a witness complex construction to study the coverage of cooling centers. One standard technique for studying the risk of heat-related mortality for a geographic area is a heat vulnerability index (HVIs) based on demographic information. We test our topological approach and an HVI on four locations (central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL) and use death times of connected components and cycles to identify most at risk regions. PH and HVI identify different locations as vulnerable, thus showing measures of coverage need to extend beyond a simple statistic. Using persistent homology along side the HVI score identifies a complementary set of regions at risk, thus the combination of the two provide a more holistic understanding of coverage.
keywords:
@@ Abstract
1 Introduction
@@
The Environmental Protection Agency has identified extreme heat as the modern leading cause of weather-related mortality in the United States [11]. The problem of heat-related mortality is expected to remain a pressing concern, especially if proactive measures are not adopted. Meehl and Tebaldi’s global coupled climate model showed that heat waves in North America are projected to intensify, occur more frequently, and persist longer in the latter half of the 21st century [25]. This escalation in heat intensity is linked to rising levels of greenhouse gases. There is evidence that the warmest day of the year will increase by 4-6°C (7.2-10.8°F) by the end of the century [29]. Heat accumulation is expected to be greater in urban areas due to human activity and construction, a phenomena coined “the urban heat island effect” [34]. The increase in temperature presents a major health hazard.
@@
Exposure to extreme heat can result in dire health outcomes including heat stroke, the exasperation of existing medical conditions, permanent neurological damage, and, in the worst cases, loss of life [3]. In particular, children and seniors are at a greater risk as they have increased difficulty adapting to heat [16]. Therefore, there is a critical need to develop tools to help locate populations most vulnerable to heat-related mortality. Often, the term“excess deaths” is used when reporting the quantity of heat-related fatalities as they are often preventable. For instance, a single heat-wave event in Europe resulted in approximately 22,000-45,000 excess deaths in 2003 [3].
@@
In a study that aimed to identify protective measures against heat-wave related deaths, Bouchama found that visiting air conditioned environments was strongly associated with better health outcomes [3]. In fact, the meta-analysis showed that those who visited air conditioned places had an approximately 66% lower likelihood of heat-related mortality. However, approximately 28.6% of housing units in the United States do not have air conditioning [24]. Furthermore, certain populations are more likely to lack air conditioning due to limited financial and social resources [24]. Therefore, it is important for cities to offer free, accessible, public cooling spaces during extreme heat events. Cooling centers can include buildings such as libraries, community centers, and senior centers [24]. It is imperative not only to increase the quantity of cooling centers but also to strategically position them in areas of highest necessity, taking into account factors such as the demographics within the city and existing cooling center locations. Currently, as Kim addresses, the distribution of cooling centers is not always optimized to maximize access [24]. Altogether, these findings motivate our emphasis on studying the geographic distribution of cooling centers within cities in the United States.
1.1 Related Work and Contributions
@@
Currently, a variety of computational tools exist to evaluate a region’s cooling center coverage. One technique to evaluate coverage is the catchment area method which involves selecting a cut-off distance and evaluating who is within that distance of a given cooling center [24, 27]. However, this approach alone overlooks important demographics that may helpful in determining areas of greatest risk. For instance, a neighborhood of primarily senior citizens with little tree canopy may need to be treated differently than other neighborhoods. To account for this, many studies compute a social vulnerability index for a given geographic area based on demographic variables that influence an individual’s ability to overcome environmental hazards [8]. In the context of extreme heat risk, they are often called heat vulnerability indices (HVIs). There are multiple methods to calculate a region’s HVI [4]. Some previous studies utilize a combination of HVI maps and geospatial statistics to inform a maximal coverage location problem, that is, finding the optimal locations for new cooling centers. For example, the GIS Network Analyst tool, which optimizes road networks, has been used in conjunction with HVI data to recommend optimal locations for future cooling centers [4, 15]. This approach, however, lacks interpretability since one of the algorithms used has some built-in heuristics that are not transparent to the user.
@@
A tool from topological data analysis (TDA), persistent homology (PH), provides a complementary perspective to these existing techniques. Firstly, TDA can evaluate coverage across a range of parameter values which bypasses the arbitrary selection of a cut-off distance. Secondly, many traditional methods, such as those that rely solely on the information provided by an HVI map, treat each census block as an isolated “island” when assessing the risk of its residents. For example, a single census block without cooling centers within its borders might be deemed highly at risk using conventional methods. In contrast, PH is able to incorporate spatial information by considering cooling centers in neighboring census blocks. Therefore, PH can analyze the surrounding areas and determine that the risk may be lower than it appears due to the abundant cooling centers in a neighboring block, for example. Lastly, our methodology considers cooling centers not only within city limits, but also in neighboring cities. In reality, during an extreme heat event, individuals who live on the boundary of a given city have the option of traveling to neighboring cities to find shelter.
@@
Persistent homology has shown success in many applications to spatial data [19, 12, 20, 13, 7]. Our method is inspired by that of Hickok et al. [19] which uses TDA to infer holes in voting site coverage. However we use a different construction of the mathematical representation of space. In particular, we use a witness complex while the authors of Hickok et al. [19] use a weighted Vietoris–Rips complex.
1.2 Organization of this paper
@@
In Section 2 we introduce the relevant background information on persistent homology. In Section 3, we present two methodologies, one based on persistent homology and the other an HVI for comparison. We present and analyze the results in Section 4 and conclude with discussions of implications and limitations of our work in Section 5
2 Background
@@
Here we present a brief overview of topological data analysis and, in particular, persistent homology. We first begin with some preliminary definitions in Sec. 2.1, followed by an introduction to persistent homology in Sec. 2.2. We point the interested reader to [9] for a more formal treatment of this material.
2.1 Definitions
@@
A simplicial complex is a shape constructed from lower dimensional building blocks such as vertices (00-simplices), edges ($1$-simplices), triangles ($2$-simplices), and higher dimensional analogs. In general, a $k$-simplex $\sigma$ is the convex hull of $k+1$ linearly independent points. A face $\tau$ of a simplex $\sigma$ is the convex hull of a subset of the points in $\sigma$; if $\tau$ is a face of $\sigma$ we write $\tau\subseteq\sigma$. A simplicial complex $\mathcal{K}$ is a set of simplices that satisfy two conditions: (1) $\mathcal{K}$ is closed under taking faces (i.e., if $\sigma\in\mathcal{K}$ and $\tau\subseteq\sigma$ then $\tau\in\mathcal{K}$) and (2) $\mathcal{K}$ is closed under intersections (i.e., if $\sigma_{1},\sigma_{2}\in\mathcal{K}$ then $\sigma_{1}\cap\sigma_{2}\in\mathcal{K}$). A filtered simplicial complex is a simplicial complex $\mathcal{K}$ and a function $f:\mathcal{K}\to\mathbb{R}$ such that $f(\tau)\leq f(\sigma)$ for $\tau\subseteq\sigma$.
From a filtered simplicial complex, one can build a nested sequence of simplicial complexes
@@
The first step in computing the “shape” of data is to first represent the data as a (filtered) simplicial complex. This can be done in numerous ways including common approaches such as a Čech or Vietoris–Rips complex [9]. Here we use a witness complex because it allows for the incorporation of two sets of points, the landmarks $L$ and the witnesses $W$ [32]. Often the landmarks are chosen as a subset of the witnesses so the resulting complex has fewer simplices than a Vietoris–Rips complex. However, this is not a requirement. The landmarks form the vertices of the simplicial complex and witnesses are used to determine which higher dimensional simplices are present. The landmarks form the 00-simplices and a $k$-simplex for $k>0$ is in the complex if it is “witnessed” by a point in $W$. This “witness” relation can be defined in different ways to create variants such as a weak or strong witness complex [32, 1, 31]. We define a filtered witness complex as a witness complex paired with a filter function $f:\mathcal{W}\to\mathbb{R}$ where $f(\sigma)$ is the distance from $\sigma$ to its witness. From this, a filtration can be built by following the same construction as defined in Eqn. 1. We define our construction of the witness complex in detail in Section 3.1.
2.2 Persistent Homology
@@
Homology is an invariant from algebraic topology used to quantify holes in different dimensions. A 0-dimensional (0D) hole represents a connected component and a 1-dimensional (1D) hole represents a cycle or a loop. Given a simplicial complex, one can compute homology; given a filtration, one can compute persistent homology. At each step in the filtration, connected components may appear and merge, and holes may form and fill in. A component or a hole is said to be born in the first step of the filtration in which it appears. A component is said to have died when it merges with another connected component, while a hole dies when it is filled in. More formally, the birth time of a homology class is $\alpha_{b}$ if the the homology class appears in $\mathcal{K}_{\alpha_{b}}$ but not $\mathcal{K}_{\alpha_{i}}$ for $i<b$. Similarly, the death time of a homology class is $\alpha_{d}$ for $d>b$ if the homology class is present in $\mathcal{K}_{\alpha_{d-1}}$ but not in $\mathcal{K}_{\alpha_{d}}$. When two components merge, the older (i.e. the one born earlier) persists while the younger dies; this is known as the elder rule [10]. The birth simplex is the first simplex that creates a homology class, while the death simplex is the last simplex added that kills the homology class.
@@
Figure 1: An example of a distribution of 11 census tract centroids (landmarks) and randomly generated cooling centers (witnesses). Note: this image was created based on a small subset of Boston, MA census tracts. Our results on Boston, MA as presented in Section 4 use census block centroids (landmarks) and OpenStreetMap cooling centers (witnesses).
3 Methods
In this section we present two methodologies for evaluating cooling center coverage. Section 3.1 introduces our topological approach using persistent homology of a filtered witness complex to identify the vulnerable locations. Subsequently, Section 3.2 defines a heat vulnerability index (HVI), inspired by existing literature. We use this more standard approach to studying vulnerability as a comparison to the topological approach. Note that our topological approach uses census blocks, however due to the availability of demographic data, the HVI score can only be computed for census tracts.
3.1 Our Construction of the Witness Complex
@@
[…]
where $d(\ell,w)$ is the distance between landmark $\ell$ and witness $w$.
Intuitively, this is a simplicial complex constructed with landmarks as the vertices (all of which are born at $\alpha=0$), and simplices built from the sets of vertices contained within distance $\alpha$ from the same witness. Note that this construction is a clique complex; that is, if all faces of a simplex are in the complex, the face is included as well. Given an increasing set of $\alpha$ values
@@
$\mathcal{W}(L,W)_{\alpha_{0}}\subseteq\mathcal{W}(L,W)_{\alpha_{1}}\subseteq\mathcal{W}(L,W)_{\alpha_{2}}\subseteq\cdots\subseteq\mathcal{W}(L,W)_{\alpha_{n}}$
@@
Figure 2: Example of select steps in the filtration. As shown in Figure 1, light green points represent the centroids (landmarks) of the geographic region and dark green diamonds (witnesses) represent the cooling center locations. All vertices (landmarks) appear in (a). (b)-(d) show the simplicial complex (in black) at various stages of the filtration. Edges are drawn between vertices when they are within a distance $\alpha$ of the same witness. Green discs show the radii at that filtration value. Note that the witnesses are draw for easier visualization but they are not vertices in the simplicial complex.
@@
For this application, we use the centroids of census blocks as the landmarks, and the locations of cooling centers as the witnesses, refer to Figure 1 as an example. This means, in the filtered witness complex, an edge is added between two landmarks when the corresponding census blocks are both within distance $\alpha$ of the same cooling center. See Figure 2 as an example of how components evolve over select steps in a filtration. When two landmarks are connected by an edge, in other words, when they have access to the same cooling center, we interpret this as the two census blocks having “equal coverage.”
Given that each city is in a different region of the country, computing distances like a Euclidean distance on a projected map of the US may distort the distances in some cities more than others. To avoid the distortion of projected maps, we compute distances between latitude and longitude points using the shortest arc length or geodesic distance. We believe this choice to be justifiable because some studies have shown straight-line measurements are an acceptable proxy to travel distances in urban areas [2, 14, 22]. Furthermore, in non-emergency situations, Boscoe et al. concluded that the difference between straight-line and driving estimates were inconsequential and, because of this, straight-line estimates are often used [14, 27]. This measure of distance is then converted into units of kilometers for easier interpretation.

### v2

# Evaluating Cooling-Center Coverage Using Persistent Homology of a Filtered Witness Complex

@@
Abstract.
In light of the increase in frequency of extreme heat events, there is a critical need to develop tools to identify geographic locations that are at risk of heat-related mortality. This paper aims to identify locations by assessing holes in cooling-center coverage using persistent homology (PH), a method from topological data analysis (TDA). Persistent homology has shown promising results in identifying holes in coverage of specific resources. We adapt these methods using a witness complex construction to study the coverage of cooling centers. We test our approach on four locations (central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL) and use death times, a measurement of the size and scale of the gap in coverage, to identify most at risk regions. For comparison, we implement a standard technique for studying the risk of heat-related mortality called a heat vulnerability index (HVI). The HVI is a numerical score calculated for a geographic area based on demographic information. PH and the HVI identify different locations as vulnerable, thus indicating a potential value of assessing vulnerability from multiple perspectives. By using the regions identified by both persistent homology and the HVI, we provide a more holistic understanding of coverage.
Key words and phrases:
@@ Abstract.
Program in Applied & Computational Mathematics, Princeton University, Princeton, NJ (erinoneil@princeton.edu).
Department of Mathematics and Computer Science, College of the Holy Cross, Worcester, MA (stymochko@holycross.edu).
1. Introduction
@@
The Environmental Protection Agency has identified extreme heat as the modern leading cause of weather-related mortality in the United States [14]. The problem of heat-related mortality is expected to remain a pressing concern, especially if proactive measures are not adopted. Meehl and Tebaldi’s global coupled climate model [30] showed that heat waves in North America are projected to intensify, occur more frequently, and persist longer in the latter half of the 21st century compared to the latter half of the 20th century. This escalation in heat intensity is linked to rising levels of greenhouse gases. There is evidence that the warmest day of the year will increase by 4-6°C (7.2-10.8°F) by the end of the century [34]. Heat accumulation is expected to be greater in urban areas due to human activity and construction, a phenomena coined “the urban heat island effect” [41].
@@
Exposure to extreme heat can result in dire health outcomes including heat stroke, the exasperation of existing medical conditions, permanent neurological damage, and, in the worst cases, loss of life [5]. Often, the term“excess deaths” is used when reporting the quantity of heat-related fatalities as they are often preventable. For instance, a single heat-wave event in Europe resulted in approximately 22,000-45,000 excess deaths in 2003 [5]. There is a critical need to develop tools to help locate vulnerable populations and identify protective measures against these heat-related deaths.
@@
In a study that aimed to identify such protective measures, Bouchama found that visiting air conditioned environments was strongly associated with better health outcomes [5]. In fact, the meta-analysis showed that those who visited air conditioned places had an approximately 66% lower likelihood of heat-related mortality. However, approximately 28.6% of housing units in the United States do not have air conditioning [28]. Furthermore, certain populations are more likely to lack air conditioning due to limited financial and social resources [28]. Therefore, it is important for cities to offer free, accessible, public cooling centers, such as libraries, community centers, and senior centers [28] during extreme heat events. It is imperative not only to increase the quantity of cooling centers but also to strategically position them in areas of highest necessity, taking into account factors such as the demographics within the city and existing cooling center locations. Recent work shows that the distribution of cooling centers is not always optimized to maximize access [28].
1.1. Related Work and Contributions
@@
Our goal is to study the coverage of cooling centers throughout a city and identify any holes in coverage. Currently, a variety of computational tools exist to evaluate a region’s cooling center coverage. One technique to evaluate coverage is the catchment area method which involves selecting a cut-off distance and evaluating who is within that distance of a given cooling center [28, 32]. However, this approach alone overlooks important demographics that may be helpful in determining areas of greatest risk. For instance, a neighborhood of primarily senior citizens with little tree canopy may need to be treated differently than other neighborhoods. To account for this, many studies compute a social vulnerability index for a given geographic area based on demographic variables that influence an individual’s ability to overcome environmental hazards [10]. In the context of extreme heat risk, they are often called heat vulnerability indices (HVIs).
There are multiple methods to calculate an HVI for a given region; see [2, 29] for a review of these approaches. Some previous studies utilize a combination of HVI maps and geospatial statistics to inform a maximal coverage location problem for finding “optimal” locations for new cooling centers. For example, work by Bradford et al. [6] and Fraser et al. [18] use an optimization approach that is informed by the HVI data. These methods, however, lack interpretability as they use proprietary ArcGIS tools. Further, they rely on choosing a fixed distance threshold like the catchment area method.
@@
A tool from topological data analysis (TDA), persistent homology (PH), provides a complementary perspective to these existing techniques. Firstly, TDA can evaluate coverage across a range of parameter values which bypasses the arbitrary selection of a cut-off distance. Secondly, many traditional methods, such as those that rely solely on the information provided by an HVI, treat each census block as an isolated “island” when assessing the risk of its residents. In contrast, PH is able to incorporate spatial information by considering cooling centers in neighboring census blocks. Lastly, our methodology considers cooling centers not only within city limits, but also in neighboring cities. [^1] See Appendix A.2 for a detailed explanation of the data and how it was collected. In reality, individuals who live on the boundary of a given city may reside closer to cooling centers in neighboring cities than in their own. We therefore allow for the assumption that in an extreme heat event, individuals may visit neighboring cities to find shelter.
@@
Persistent homology has shown success in many applications to spatial data [15, 24, 16, 9, 11]. Our method is inspired by that of Hickok et al. [23] which uses TDA to infer holes in voting site coverage. However, we use a different construction of the mathematical representation of space.
1.2. Organization of this paper
@@
In Section 2 we introduce the relevant background from TDA and introduce on persistent homology. In Section 3, we present our persistent homology-based method as well as the HVI we use for comparison. We present and analyze the results in Section 4 and conclude with discussions of implications and limitations of our work in Section 5.
2. Background
@@
Here we present a brief overview of topological data analysis and, in particular, persistent homology. We first begin with some preliminary definitions in Section 2.1, followed by an introduction to persistent homology in Section 2.2. We point the interested reader to [12] for a more formal treatment of this material.
2.1. Definitions
@@
A simplicial complex is a shape constructed from lower dimensional building blocks such as vertices (0-simplices), edges ($1$-simplices), triangles ($2$-simplices), and higher dimensional analogs. In general, a $k$-simplex $\sigma$ is the convex hull of $k+1$ linearly independent points. A face $\tau$ of a simplex $\sigma$ is the convex hull of a subset of the points in $\sigma$; if $\tau$ is a face of $\sigma$, we write $\tau\subseteq\sigma$. A $k$-simplex $\sigma$ can be identified uniquely by the set of 0-simplices $\tau$ such that $\tau\subseteq\sigma$. A simplicial complex $\mathcal{K}$ is a set of simplices that satisfy two conditions: (1) $\mathcal{K}$ is closed under taking faces (i.e., if $\sigma\in\mathcal{K}$ and $\tau\subseteq\sigma$, then $\tau\in\mathcal{K}$) and (2) $\mathcal{K}$ is closed under intersections (i.e., if $\sigma_{1},\sigma_{2}\in\mathcal{K}$, then $\sigma_{1}\cap\sigma_{2}\in\mathcal{K}$). A filtered simplicial complex is a simplicial complex $\mathcal{K}$ and a function $f:\mathcal{K}\to\mathbb{R}$ such that $f(\tau)\leq f(\sigma)$ for $\tau\subseteq\sigma$. From a filtered simplicial complex, one can build a nested sequence of simplicial complexes
@@
2.1.1. Constructing a Simplicial Complex from Data
The first step in computing the “shape” of data is to first represent the data as a (filtered) simplicial complex. This can be done in numerous ways, two of the most common being the Čech or Vietoris–Rips complexes [12]. Given a collection of points $X=\{x_{1},\ldots,x_{n}\}$ in a metric space $(M,d)$, the Čech complex at filtration parameter $\alpha$ is the simplicial complex with vertices $X$ and simplices $\{x_{i_{0}},\ldots x_{i_{k}}\}$ if $\bigcap_{j}B(x_{i_{j}},\alpha)$ is nonempty, where $B(x,\alpha)=\{y\in M:d(x,y)\leq\alpha\}$. The Čech complex is topologically equivalent to the nerve of the union of balls, $\bigcup_{j}B(x_{i_{j}},\alpha)$ by the Nerve theorem [4] so long as the balls are convex. The Čech complex can be used to create a filtration by using an increasing sequence of filtration parameter values, $\alpha_{0}<\alpha_{1}<\ldots<\alpha_{N}$.
The Čech complex is very useful in theory, but in practice is very computationally prohibitive. It is common to instead use the Vietoris–Rips (VR) complex as it is an approximation of the Čech complex and is significantly faster to compute. [^2] Computing the Čech complex requires knowledge of the metric space in order to check if the intersection of $k$ balls is nonempty. In comparison the VR complex requires only the pairwise distances between all points. The VR complex at filtration parameter $\alpha$ is the simplicial complex with vertices $X$ and simplices $\{x_{i_{0}},\ldots x_{i_{k}}\}$ if $d(x_{i_{j}},x_{i_{\ell}})<2\alpha$ for all $j$ and $\ell$. Just as with the Čech complex, the VR complex gives rise to a filtration when computed for an increasing sequence of filtration parameter values.
@@ 2.1.1. Constructing a Simplicial Complex from Data
Here we use a different approach for creating a simplicial complex from data. In particular, we use the witness complex $\mathcal{W}$ because it allows for the incorporation of two sets of points, the landmarks $L$ and the witnesses $W$ [38]. Often the landmarks are chosen as a subset of the witnesses, however, this is not a requirement. [^3] The VR complex remains computationally prohibitive for large point sets. The witness complex can capture similar topological structure in a way that is more computationally tractable. In this setting, the full point cloud serves as the set of witnesses while a subset of the points serve as the landmarks. The resulting complex has fewer simplices than a Vietoris–Rips complex, thus speeding up computations. This, however, is not our motivation for using the witness complex. In our case, the landmarks and witnesses will come from two separate sets of points; the landmarks represent neighborhoods throughout a city and the witnesses represent the locations of the cooling centers. The landmarks form the vertices of the simplicial complex and witnesses are used to determine which higher dimensional simplices are present. That is, a $k$-simplex for $k>0$ is in the complex if it is “witnessed” by a point in $W$. This “witness” relation can be defined in different ways to create variants such as a weak or strong witness complex [38, 1, 36]. We define a filtered witness complex as a witness complex paired with a filter function $f:\mathcal{W}\to\mathbb{R}$ where $f(\sigma)$ is the distance from $\sigma$ to the closest witness. This gives rise to a filtration using the same definition as in Eqn. 1. We define our construction of the witness complex in Section 3.1.
2.2. Persistent Homology
@@
Homology is an invariant from algebraic topology used to quantify holes in different dimensions. A 0-dimensional (0D) hole represents a connected component and a 1-dimensional (1D) hole represents a cycle or a loop. Given a simplicial complex, one can compute homology; given a filtration, one can compute persistent homology. At each step in the filtration, connected components may appear and merge, and holes may form and fill in. A component or a hole is said to be born in the first step of the filtration in which it appears. A component is said to have died when it merges with another connected component, while a hole dies when it is filled in. More formally, the birth time of a homology class is $\alpha_{b}$ if the the homology class appears in $\mathcal{K}_{\alpha_{b}}$ but not $\mathcal{K}_{\alpha_{i}}$ for $i<b$. Similarly, the death time of a homology class is $\alpha_{d}$ for $d>b$ if the homology class is present in $\mathcal{K}_{\alpha_{d-1}}$ but not in $\mathcal{K}_{\alpha_{d}}$. When two components merge, the older (i.e. the one born earlier) persists while the younger dies; this is known as the elder rule [13].
The birth simplex is the first simplex to appear in the filtration that creates a homology class, while the death simplex is the last simplex added that kills the homology class. Note that in general, a death simplex may not be unique; for example, if three vertices are all born at time 0 and the three pairwise edges are all added at the same filtration value, all three edges could be considered the death simplex. In real datasets, this is rarely an issue: distances are typically distinct, so edges and higher-dimensional simplices enter at unique filtration values.
@@
Figure 1. An example of $1$ census tract centroids (landmarks) and 7 cooling centers (witnesses). Note: this image was created based on a small subset of Boston, MA census tracts and the witness locations were generated randomly. To see the full city of Boston with the true locations of cooling centers, see Figure 11.
3. Methods
In this section we present two methodologies for evaluating cooling center coverage. Section 3.1 introduces our topological approach using persistent homology of a filtered witness complex to identify the vulnerable locations. Section 3.2 defines a heat vulnerability index (HVI), inspired by existing literature, to serve as a comparison to the topological approach.
In order to represent the spatial layout of the city, we use geographic units that are used by the U.S. Census Bureau to aggregate data. Our topological approach uses census blocks, the smallest geographic unit used by the Census Bureau, while the HVI was computed at the census tract level. This is because the demographic data needed to compute the HVI is only available at the census tract level. See Appendix A for more details.
3.1. Our Construction of the Witness Complex
@@
where $d(\ell,w)$ is the distance between landmark $\ell$ and witness $w$. Note that equation above includes all landmark points as 0-simplices for all scales $\alpha\geq 0$.
Intuitively, this is a simplicial complex constructed with landmarks as the vertices (all of which are born at $\alpha=0$), and simplices built from the sets of vertices contained within distance $\alpha$ from the same witness. Note that this construction is a clique complex; that is, if all faces of a simplex are in the complex, the face is included as well. This makes computing the witness complex similar in computation time to the VR complex since all that is needed are the pairwise distances between all pairs of landmark and witness points. Given an increasing sequence of $\alpha$ values
@@
$\mathcal{W}(L,W)_{0}\subseteq\mathcal{W}(L,W)_{\alpha_{0}}\subseteq\mathcal{W}(L,W)_{\alpha_{1}}\subseteq\cdots\subseteq\mathcal{W}(L,W)_{\alpha_{n}}$
@@
Figure 2. Example of select steps in the filtration. As shown in Figure 1, light green points represent the centroids (landmarks) of the geographic region and dark green diamonds (witnesses) represent the cooling center locations. All vertices (landmarks) appear in (a). (b)-(d) show the simplicial complex (in black) at various stages of the filtration. Edges are drawn between vertices when they are within a distance $\alpha$ of the same witness. Green discs show the radii at that filtration value. Note that the witnesses are drawn for easier visualization but they are not vertices in the simplicial complex.
@@
In this work, the centroids of census blocks are used as the landmarks, and the locations of cooling centers are used as the witnesses. See Figure 1 for an example. This means, in the filtered witness complex, an edge is added between two landmarks when the corresponding census blocks are both within distance $\alpha$ of the same cooling center. See Figure 2 as an example of how components evolve over select steps in a filtration. When two landmarks are connected by an edge, in other words, when they have access to the same cooling center, we interpret this as the two census blocks having “equal coverage.”
3.1.1. Interpretation and Visualization
In order to further identify and visualize regions of poor coverage from our topological approach, we use the death simplices as defined in Section 2.2.444Another approach to identifying “holes” in coverage from 1-dimensional homology is using cycle representatives. However, because cycle representatives are not unique and finding “optimal” cycles is an open and ongoing area of research, we leave this as potential future work. Death simplices for 0-dimensional homology are edges (1-simplices) which represent the edge that is added to connect two previously distinct components. These edges connect components that previously had differing levels of cooling center coverage. For $1$-dimensional homology, death simplices are triangles (2-simplices) which represent the final part of the hole to be covered. These regions are the convex hull of three census blocks between which there is a lack of cooling centers. A death simplex with a higher death time indicates worse coverage in that region.
As discussed previously, death simplices need not be unique in theory. However in this case, because simplices of dimension 1 or greater are added based on distances calculated between geographical points, then we can guarantee that at most one simplex is added at any given filtration value and thus the death simplices are unique. [^5] When identifying the 0-dimensional death simplices, we recompute the persistence points after slightly perturbing the birth times of the 0-simplices. This ensures that each 0-dimensional feature can be uniquely matched to the corresponding death simplex. The perturbations are carefully chosen so as not to impact the order in which simplices of dimension 1 or greater are added, thus not changing the death times of persistence points nor the corresponding death simplices.